A beam of protons with a velocity 4 x 105 m/sec enters a uniform magnetic field of 0.3 tesla at an angle of 60o to the magnetic field. Find the radius of the helical path taken by the proton beam. Also find the pitch of the helix (which is the distance travelled by a proton in the beam parallel to the magnetic field during one period of rotation). 

ans:

radius =0.012m

pitch=0.044m


Grade:10

1 Answers

ROSHAN MUJEEB
3 years ago
dear student,
The radius of helical path:r=Bqmv⊥​​
=Bqmvsinθ​=(0.3)(1.6×10−19)(1.67×10−27)(4×105)(sin60o)​
=1.2×10−2m

Pitch of helical path,p=v∥​×T
P=(Bq2πm​)(vcosθ)
=(03.)(1.6×10−19)(2π)(1.67×10−27)(4×105)(cos60o)​=4.37×10−2m
hope it helps